### v1

# A Common Origin of Asymmetric Self-interacting Dark Matter and Dirac Leptogenesis

Abstract
Assuming dark matter to be asymmetric as well as self-interacting and neutrinos to be Dirac fermions, we propose a framework to address the observed baryon imbalance of the universe. We add three right-handed neutrinos $\nu_{R_{i}},\,{i=1,2,3}$, one singlet fermion $\chi$, a doublet fermion $\psi$, and heavy scalar doublets $\eta_{i},\,{i=1,2}$ to the Standard Model. Both $\chi$ and $\psi$ are fermions with non-zero charge under an extended $U(1)_{B-L}\times U(1)_{D}$ symmetry. Additionally, a $\mathcal{Z}_{2}$ symmetry is imposed, where the singlets $\chi$, $\nu_{R}$, and $\eta$ are negative and the doublet $\psi$ is positive. Given the assumption that neutrinos are Dirac particles, $B-L$ turns into an exact symmetry of the universe. The CP-violating out-of-equilibrium decay of heavy scalar $\eta$ generates an equal and opposite $B-L$ asymmetry among the left-handed ($\nu_{L}$) and right-handed ($\nu_{R}$) neutrinos. The $\nu_{L}-\nu_{R}$ equilibration process does not take place until below the Electroweak phase transition scale because of tiny Yukawa couplings. During this time, Sphaleron processes, which are active at temperatures higher than 100 GeV, transform a portion of the $B-L$ asymmetry stored in left-handed neutrinos into baryon asymmetry. MeV scale gauge boson $Z^{\prime}$ of $U(1)_{D}$ sector mediates both annihilation of symmetric dark matter component and self-interaction among dark matter particles. Moreover, $Z^{\prime}$ mixes with the Standard Model Z-boson and provides a portal for dark matter direct detection.
I Introduction
The Standard Model (SM) is a remarkably successful theory explaining interactions of elementary particles. Yet several mysteries remain unanswered in SM, of which the non-zero neutrino mass, the observed baryon asymmetry of the Universe, and the candidature problem of dark matter (DM), etc., are the most prominent issues. The rotation velocities of stars in galaxies, gravitational lensing, anisotropies in the CMB, etc., hint the existence of an electromagnetically inert form of matter. This new form of matter is named dark matter (DM). The particle nature of DM is yet to be revealed. WMAP Hinshaw:2012aka and PLANCK Aghanim:2018eyx data (at 68% CL) quantify the relic baryon and DM density as: $\Omega_{\rm B}h^{2}=0.02237\pm 0.00015$, $\Omega_{\rm DM}h^{2}=0.1200\pm 0.0012$; where $\Omega_{\rm DM}$ is the ratio of DM density to the critical density of the Universe and $h=\text{Hubble Parameter}/(100\;\text{km}~{}\text{s}^{-1}\text{Mpc}^{-1})$ is the reduced Hubble constant. Observations also indicate that the present Universe is highly baryon asymmetric. The observed Baryon Asymmetry of the Universe (BAU) can be quantified as Aghanim:2018eyx ,
$\eta_{B}=\frac{n_{B}-n_{\overline{B}}}{n_{\gamma}}\simeq 6.2\times 10^{-10},$ (1)
where, $n_{B}$, $n_{\overline{B}}$ and $n_{\gamma}$ are the number densities of baryon, antibaryon and photon respectively. Based on Sakharov conditions Sakharov:1967dj ; sakharov.67 , the observed baryon asymmetry has been addressed in frameworks of leptogenesis fukugita.86 ; baryo_lepto_group_1 ; baryo_lepto_group_2 ; baryo_lepto_group_3 ; baryo_lepto_group_4 ; baryo_lepto_group_5 ; baryo_lepto_group_6 ; baryo_lepto_group_7 ; baryo_lepto_group_8 ; baryo_lepto_group_9 ; Abada:2006ea ; Pilaftsis:2003gt ; Pilaftsis:2005rv ; Kaplan:2009ag ; Feng:2012jn ; Kolb:1990vq ; Ibe:2011hq ; Feng:2013wn ; Narendra:2018vfw . WMAP and Planck data imply that $\Omega_{\rm DM}\approx 5\Omega_{\rm B}$. It triggers to assume DM to have a similar origin as that of baryons and address both phenomenologies in a common framework old_DM_Baryon_asy_1 ; old_DM_Baryon_asy_2 ; old_DM_Baryon_asy_3 ; old_DM_Baryon_asy_4 ; old_DM_Baryon_asy_5 ; old_DM_Baryon_asy_6 ; old_DM_Baryon_asy_7 ; old_DM_Baryon_asy_8 ; Asydm_models1_1 ; Asydm_models1_2 ; Asydm_models1_3 ; Asydm_models1_4 ; Asydm_models1_5 ; Asydm_models1_6 ; Asydm_models1_7 ; Asydm_models1_8 ; Asydm_models1_9 ; Asydm_models1_10 ; Asydm_models2_1 ; Asydm_models2_2 ; Asydm_models2_3 ; Asydm_models2_4 ; Asydm_models2_5 ; Asydm_models2_6 ; Asydm_models2_7 ; Asydm_models2_8 ; Asydm_models2_9 ; Arina:2011cu ; Arina:2012fb ; Arina:2012aj ; Narendra:2017uxl ; Asydm_review_1 ; Asydm_review_2 .
From cosmological standpoint, DM is assumed to be a cold, collisionless fluid by the ${\rm\Lambda CDM}$ model. ${\rm\Lambda CDM}$ model has proved incredibly successful in explaining the large-scale structures of the Universe. Nevertheless, a number of discrepancies have surfaced at smaller scales, such as the core-cusp problem, the missing satellite problem, and the too-big-to-fail problem etc Tulin:2017ara ; Bullock:2017xww . These discrepancies are more noticeable at the dwarf galaxy scale and progressively diminish to the cluster scale, which corresponds with $\Lambda{\rm CDM}$ predictions at large scales. In order to mitigate these anomalies, self-interacting dark matter (SIDM) Carlson:1992fn ; deLaix:1995vi ; Spergel:1999mh was suggested as a substitute for traditional cold dark matter. A velocity-dependent self-scattering of order $\sigma/m\sim 1\;{\rm cm}^{2}/{\rm g}\approx 2\times 10^{-24}\;{\rm cm}^{2}/{\rm GeV}$ has been demonstrated to alleviate these anomalies Buckley:2009in ; Feng:2009hw ; Feng:2009mn ; Loeb:2010gj ; Zavala:2012us ; Vogelsberger:2012ku ; Bringmann:2016din ; Kaplinghat:2015aga ; Kamada:2020buc ; Aarssen:2012fx ; Tulin:2013teo . DM models with a light mediator is well suited to realise such velocity dependence Kouvaris:2014uoa ; Bernal:2015ova ; Kainulainen:2015sva ; Hambye:2019tjt ; Cirelli:2016rnw ; Kahlhoefer:2017umn ; Ghosh:2021wrk . However a common issue with SIDM is its under-abundant relic due to efficient annihilation into the light mediator Dutta:2021wbn ; Borah:2021yek ; Borah:2021pet ; Borah:2021rbx ; Dutta:2024ydz . However, if SIDM is assumed to have similar origin as that of baryons, correct relic density can be achieved Dutta:2022knf ; Borah:2024wos .
Because there are no right-handed neutrino fields, neutrinos have no mass within the SM framework. However, neutrino flavor oscillation Abe:2011sj ; Abe:2013hdq ; Abe:2011fz ; DoubleChooz:2019qbj ; An:2012eh ; An:2017osx ; Adey:2018zwh ; Ahn:2012nd ; Adamson:2013whj ; Adamson:2014vgd suggests that the neutrino mass is non zero but small. Beyond the SM framework invoking lepton number violating dimension-5 operator $LLHH/\Lambda$ Weinberg:1979sa is a lucrative method of explaining the small mass of Majorana neutrinos. The scale of new physics is $\Lambda$, whereas the lepton and Higgs doublets are denoted by $L,H$. In this way, neutrinos obtain a Majorana mass of order $m_{\nu}=\langle H\rangle^{2}/\Lambda$ during electroweak phase transition (EWPT). This novel framework is referred to as seesaw mechanism includes type-I Minkowski:1977sc ; Yanagida:1981xy ; GellMann:1980vs ; Mohapatra:1979ia , type-II type_II_seesaw_1 ; type_II_seesaw_2 ; type_II_seesaw_3 ; type_II_seesaw_4 ; type_II_seesaw_5 ; type_II_seesaw_6 ; type_II_seesaw_7 , type-III type_III_seesaw or their variants ray_volkas_review . This mechanism works well for $\Lambda\sim{\cal O}(10^{14}){\rm GeV}$ for neutrino masses below eV. Neutrinoless double beta decay experiments indicative of the Majorana nature of neutrinos onu_beta_expt are being investigated. However, as of yet, no corroborating results have been found. Consequently, there is still a chance that neutrinos could be Dirac fermions.
Considering that neutrinos are Dirac particles (i.e., $B-L$ is an exact symmetry), Dirac Leptogenesis can account for the observed baryon asymmetry of the Universe Dirac_leptogenesis_1 ; Dirac_leptogenesis_2 ; Dirac_leptogenesis_3 ; Dirac_leptogenesis_4 ; Dirac_leptogenesis_5 ; Dirac_leptogenesis_6 ; Dirac_leptogenesis_7 . The fundamental mechanism of Dirac Leptogenesis is that the equilibration time between left and right-handed neutrinos mediated via Yukawa interaction with SM Higgs (i.e. $y\overline{\nu_{R}}H\nu_{L}$) is much less than the $(B+L)$ violating sphaleron transitions above electroweak phase transition. Therefore, demanding $B-L=B-(L_{\rm SM}+L_{\nu_{R}})=0$Cerdeno:2006ha , a net $B-L_{\rm SM}$ can be generated in terms of $L_{\nu_{R}}$. $\nu_{R}$ being $SU(2)_{L}$ singlet, $L_{\nu_{R}}$ remains unaffected by sphalerons. The non-zero $B-L_{\rm SM}$ is, however, converted to a net $B$ asymmetry via $B+L$ violating sphaleron transitions Dick ; Buchmuller:2000as ; Cerdeno:2006ha ; Gu:2006dc ; Gu:2007mc ; Murayama:2002je ; Thomas:2005rs ; Borah:2016zbd ; Perez:2021udy .
To integrate all these avenues, in this paper, we construct a framework for explaining DM and baryon asymmetry considering Dirac neutrinos. The asymmetry in both sectors is generated by the CP-violating out-of-equilibrium decay of a heavy scalar doublet $\eta$ into $\ell\nu_{R}$ and $\psi\chi$, where $\ell$ is SM lepton, $\nu_{R}$ is right handed neutrino (RHN), and $\psi$, $\chi$ are respectively a vector-like fermion doublet and a vector-like fermion singlet that comprises the dark sector. The lepton asymmetry in the channel $\eta\to\ell\nu_{R}$ is then converted to a net baryon asymmetry via $B+L$ violating sphaleron transitions. Being a vector-like Dirac fermion, asymmetry in the channel $\eta\to\psi\chi$ remains intact Arina:2011cu ; Arina:2012fb ; Arina:2012aj ; Narendra:2017uxl . The dark sector is gauged under a $U(1)_{D}$ symmetry. The corresponding gauge boson $Z^{\prime}$ mediates DM self interaction. The lightest dark sector particle $\chi$ which is odd under $\mathcal{Z}_{2}$ acts as the SIDM candidate. The symmetric DM component annihilates away into a pair of $Z^{\prime}$, yielding a vanishingly small symmetric relic. We also impose $U(1)_{B-L}\times\mathcal{Z}_{2}$ symmetry which ensures that the neutrinos are Dirac. Meanwhile, $\mathcal{Z}_{2}$ symmetry is softly broken by a term $\mu^{2}\eta^{\dagger}H$, generating Dirac mass for neutrinos. A schematic of the framework is shown in Fig. 1.
Figure 1: Schematic of the model: A heavy scalar $\eta$ decays into visible ($l\nu_{R}$) and dark sectors ($\psi\chi$), where the dark sector includes the particles charged under $U(1)_{D}$ symmetry. The lightest dark sector particle $\chi$ is the DM candidate.
The paper is organized as follows. In sec. II, we discuss the proposed model, while in sec. III, we explain the Dirac mass of light neutrinos. Section IV is devoted to explain baryogenesis via leptogenesis from the decay of heavy particles $\eta$, while section V describes the mechanism for DM relic abundance. The phenomenology of SIDM is discussed in section VI, followed by constraints from direct search in sec. VII. We finally conclude in sec. VIII.
II The Model
We extend the SM both by new symmetries and particles. The new particles include: three RHNs $\nu_{{R}_{\alpha}}$, $\alpha=1,2,3$, a heavy scalar doublet $\eta$, and a pair of vector-like leptons, one singlet ($\chi$) and the other doublet ($\psi$) under $SU(2)$. A discrete symmetry $\mathcal{Z}_{2}$ is introduced under which $\eta,\nu_{R}$ and $\chi$ are odd, while $\psi$ and all other SM particles are even. The fermions $\psi$ and $\chi$ also carry quantum numbers under $U(1)_{D}$ symmetry. $\chi$ is the lightest of dark sector particles, hence the DM candidate. $U(1)_{D}$ symmetry is spontaneously broken and MeV scale $U(1)_{D}$ gauge boson $Z^{\prime}$ mediates DM self-interaction. The overall $B-L$ global symmetry remains intact because neutrinos are considered to be Dirac fermions. The particle content beyond those in the SM and their charge assignments are shown in Table 1.
Parameter $U(1)_{B-L}$ $U(1)_{D}$ $\mathcal{Z}_{2}$
$\eta=(\eta^{+},\eta^{0})^{T}$ 0 0 -
$\nu_{R}$1 0 -
$\psi=(\psi^{0},\psi^{-})^{T}$1 1 +
$\chi$1 1 -
Table 1: Quantum numbers of the new particles under the extended symmetry.
Heavy scalar $\eta$ undergoes CP-violating out-of-equilibrium decay. This decay creates asymmetries in both lepton and dark sectors simultaneously Arina:2011cu . The process $\eta\to\ell\nu_{R}$, creates equal and opposite lepton asymmetry in left and right-handed sectors. $B+L$ violating sphaleron transitions converts the lepton asymmetry in the left-handed sector to a net baryon asymmetry. Asymmetry in the right-handed sector remains unaffected until the temperature falls much below the electroweak phase transition. DM $\chi$ attains thermal equilibrium in early universe via processes like $\psi\psi\to\chi\chi$, mediated by $Z^{\prime}$. However, $\chi$ annihilates efficiently into pairs of $Z^{\prime}$ via t-channel and a negligible symmetric DM component is left. The correct DM relic density is obtained from the CP-violating out-of-equilibrium decay of heavy scalar $\eta$. The relevant Lagrangian is given by:
$\begin{split}\mathcal{L}&\supset\bar{\psi}i\gamma^{\mu}D_{\mu}\psi+\bar{\chi}i\gamma^{\mu}D^{\prime}_{\mu}\chi+M_{\psi}\bar{\psi}\psi+M_{\chi}\bar{\chi}\chi\\ &+\left[f_{kl}\overline{\ell_{k}}\tilde{\eta}{\nu_{R}}_{l}+\lambda\overline{\psi}\tilde{\eta}\chi+{\rm h.c.}\right]-V(H,\eta)\,,\end{split}$ (2)
where $\tilde{\eta}=i\tau_{2}\eta^{*}$, $\tau_{2}$ being the second Pauli spin matrix and,
$D_{\mu}$ $=$ $\partial_{\mu}-i\frac{g}{2}\tau^{i}W^{i}_{\mu}-i\frac{g_{{}_{Y}}}{2}YB_{\mu}-ig^{\prime}Y^{\prime}(Z^{\prime})_{\mu},$
$D^{\prime}_{\mu}$ $=$ $\partial_{\mu}-ig^{\prime}Y^{\prime}(Z^{\prime})_{\mu},$ (3)
and the scalar potential is given by,
$V(H,\eta)$ $=$ $-M^{2}_{H}H^{\dagger}H+\lambda_{H}(H^{\dagger}H)^{2}$ (4)
$+$ $M_{\eta}^{2}\eta^{\dagger}\eta+\lambda_{\eta}(\eta^{\dagger}\eta)^{2}+\lambda_{H\eta}(H^{\dagger}H)(\eta^{\dagger}\eta).$
The gauge boson $Z^{\prime}$ of the spontaneously broken $U(1)_{D}$ symmetry facilitates DM self-interaction. Hence, in this model, the observed DM relic is dominantly asymmetric. Kinetic mixing between $Z^{\prime}$ and SM $Z$ enables direct detection of DM in terrestrial laboratories. We elaborate these phenomenologies in respective sections.
III Dirac mass of neutrinos
Imposed $\mathcal{Z}_{2}$ symmetry is softly broken McDonald:2007ka ; Heeck:2013vha by the Lagrangian term:
$\mathcal{L}_{soft}=-\mu^{2}H^{\dagger}\eta+{\rm h.c.}$ (5)
Therefore, Dirac mass of the neutrinos can be generated as shown in the Fig. 2.
Figure 2: Feynman diagram responsible for generation of Dirac neutrino mass.
After integrating out the heavy field $\eta$, the Dirac mass is obtained as,
$M_{\nu}=\frac{f\langle H\rangle\mu^{2}}{M_{\eta}^{2}}\,,$ (6)
where $\langle H\rangle=174$ GeV, is the vacuum expectation value of the SM Higgs. From Eq. 6, we get neutrino mass of order $0.1$ eV, for $\frac{\mu}{M_{\eta}}\approx 10^{-4}$ and $f\sim 10^{-4}$. The smallness of $\mu$ compared to $M_{\eta}$ justifies the soft breaking.
IV Baryogenesis via Dirac Leptogenesis from $\eta$ decay
Initially the heavy scalar $\eta$ is in thermal equilibrium due to its gauge and Yukawa interactions at a temperature above its mass scale. As the Universe expands and cools down, the heavy scalar goes out-of-equilibrium and decays at a temperature $T\sim M_{\eta}$. We consider two copies of the scalar $\eta$ to achieve the CP-violation. The CP-violating out-of-equilibrium decay generates an asymmetry in both visible ($\ell\nu_{R}$) and dark sectors ($\psi\chi$). The decay width of the $\eta$ particle can be written as,
$\Gamma_{\eta}\simeq\frac{1}{8\pi}(f^{2}+\lambda^{2})M_{\eta},$ (7)
where $f$ and $\lambda$ are the Yukawa couplings. In presence of pairs of $\eta$-particles and their interactions, the diagonal mass terms for $\eta$ in Eq. 4 can be replaced by Ma:1998dx ; Arina:2011cu ,
$\frac{1}{2}\eta^{\dagger}_{a}(M_{+}^{2})_{ab}\eta_{b}+\frac{1}{2}(\eta^{*}_{a})^{\dagger}(M_{-}^{2})_{ab}\eta^{*}_{b}\,,$ (8)
Where
$M^{2}_{\pm}=\begin{pmatrix}M_{1}^{2}-iC_{11}&-iC_{12}^{\pm}\cr\\ -iC_{21}^{\pm}&M_{2}^{2}-iC_{22}\end{pmatrix},$ (9)
here $C_{ab}^{+}=\Gamma_{ab}M_{b}$, $C_{ab}^{-}=\Gamma^{*}_{ab}M_{b}$ and $C_{aa}=\Gamma_{aa}M_{a}$ with
$\Gamma_{ab}M_{b}=\frac{1}{8\pi}\left(M_{a}M_{b}\lambda_{a}\lambda_{b}^{*}+M_{a}M_{b}\mathop{\sum_{k,l}}f_{akl}^{*}f_{bkl}\right).$ (10)
Figure 3: The interference of Tree level and Self-energy correction diagrams that contribute to the CP violation.
After diagonalizing the mass matrix given in Eq. 9, we get two mass eigenstates $\xi_{1}^{\pm}$ and $\xi_{2}^{\pm}$ with two mas eigenvalues $M_{\xi_{1}}$ and $M_{\xi_{2}}$. Their decay provide CP-asymmetry via the interference of tree level and one-loop self-energy correction diagrams, shown in Fig. 3111Note that the mass eigenstates $\xi_{1}^{+}$ and $\xi_{1}^{-}$ (similarly $\xi_{2}^{+}$ and $\xi_{2}^{-}$) are not CP conjugates of each other even though they are degenerate mass eigenstates.. The generated asymmetry in the visible sector can be quantified as,
$\epsilon_{L}$ $=$ $[B_{L}(\xi_{1}^{-}\rightarrow l^{-}\nu_{R})-B_{L}(\xi_{1}^{+}\rightarrow(l^{-})^{c}\nu_{R}^{c})]$ (11)
$=$ $-\frac{Im\left(\lambda_{1}^{*}\lambda_{2}\mathop{\sum_{k,l}}f^{*}_{1kl}f_{2kl}\right)}{8\pi^{2}(M_{2}^{2}-M_{1}^{2})}\left[\frac{M_{1}^{2}M_{2}}{\Gamma_{1}}\right]\,,$
where $B_{L}$ is the branching ratio for $\xi_{1}^{\pm}\rightarrow l^{\pm}\nu_{R}$. We can estimate the generated lepton asymmetry in the visible sector using the relevant Boltzmann equations governing the evolution of the number density of $\xi_{1}$ and the lepton asymmetry:
$\begin{split}\frac{dY_{\xi_{1}}}{dx}&=-\frac{x}{H(M_{\xi_{1}})}s<\sigma|v|_{(\xi_{1}\xi_{1}\rightarrow All)}>\left[Y_{\xi_{1}}^{2}-{Y_{\xi_{1}}^{eq}}^{2}\right]\\ &-\frac{x}{H(M_{\xi_{1}})}\Gamma_{(\xi_{1}\rightarrow All)}\left[Y_{\xi_{1}}-Y_{\xi_{1}}^{eq}\right]\\ \frac{dY_{L}}{dx}&=\epsilon_{L}\frac{x}{H(M_{\xi_{1}})}\Gamma_{(\xi_{1}\rightarrow All)}B_{L}\left[Y_{\xi_{1}}-Y_{\xi_{1}}^{eq}\right]\,,\end{split}$ (12)
where the $x=M_{\xi_{1}}/T$. The lepton and $\xi$ abundance are defined as $Y_{\xi,L}\equiv n_{\xi,L}/s$, where $s=(2\pi^{2}/45)g_{*}T^{3}$ is the entropy density of the Universe.
Figure 4: The lepton asymmetry from $\xi_{1}$ decay. The Blue line shows the evolution of lepton asymmetry for $\epsilon_{L}=4.5\times 10^{-7}$. The Red line shows the evolution of $\xi_{1}$ particles. The Green(dotted) line shows the equilibrium number density of $\xi_{1}$.
Fig. 4 depicts lepton asymmetry $Y_{L}$ and comoving number density of $\xi_{1}$, i.e., $Y_{\xi_{1}}$, as a function of the dimensionless parameter $x$. Coupling constants $f$ is taken as $10^{-4}$ and $\lambda$ is taken as $10^{-7}$. The typical value of the cross-section is taken as $\sigma|v|_{(\xi_{1}\xi_{1}\rightarrow All)}=10^{-25}\,{\rm GeV}^{-2}$. The solid red and blue curves show the evolution of number densities of $\xi_{1}$ and lepton asymmetry respectively. As temperature falls below the mass of $\xi_{1}$ (i.e., $x>1$), it decouples from the thermal bath (see Fig. 4). As $\xi_{1}$ starts decaying to $\ell^{-}\nu_{R}$, lepton asymmetry develops proportional to $B_{L}$ and $\epsilon_{L}$. Once the decay of $\xi_{1}$ is complete, the lepton asymmetry settles to a constant value. In Fig. 4, we have taken the branching ratio $B_{L}=0.5$, $\epsilon_{L}$ = $5\times 10^{-7}$ and the heavy scalar mass to be $M_{\xi_{1}}$ = $10^{10}$ GeV. Electroweak Sphalerons that violate $B+L$ (but conserve $B-L$) transfer a portion of the lepton asymmetry to a net baryon asymmetry $Y_{B}=-0.55Y_{L}$.
V Asymmetric Dark matter from $\eta$ decay
At temperature above its mass scale, vector-like fermion doublet $\psi$ remains in thermal equilibrium in the early Universe. DM candidate $\chi$, in spite of being a SM gauge singlet, attains thermal equilibrium via the process $\psi\psi\to\chi\chi$ shown in Fig. 5 mediated by the light vector boson $Z^{\prime}$. This is because the dark gauge coupling $g^{\prime}$ is sufficiently large ($g^{\prime}\sim 0.1$) as required for sufficient self-interaction (discussed in Sec. VI).
Figure 5: Feynman diagram responsible for bringing DM into thermal equilibrium.
However, the same gauge coupling also facilitates efficient DM $\chi$ annihilation into vector bosons $Z^{\prime}$ through the process shown in Fig. 6.
Figure 6: Annihilation of the symmetric component of the DM.
Consequently, the relic density of the symmetric part becomes vanishingly under-abundant. Symmetric relic density can be quantified from the Boltzmann equation in terms of $Y^{sym}_{\chi}$ and $x=M_{\chi}/T$, as follows,
$\frac{dY^{sym}_{\chi}}{dx}=-\frac{s(M_{\chi})}{x^{2}H(M_{\chi})}\langle\sigma v\rangle_{\overline{\chi}\chi\to~{}Z^{\prime}Z^{\prime}}\left(Y^{2}_{\chi}-\left(Y^{eq}_{\chi}\right)^{2}\right)\,,$ (13)
where $\langle\sigma v\rangle_{\overline{\chi}\chi\to~{}Z^{\prime}Z^{\prime}}$ is the thermally averaged cross-section for $\overline{\chi}\chi\to Z^{\prime}Z^{\prime}$, which is,
$\langle\sigma v\rangle_{\overline{\chi}\chi\to~{}Z^{\prime}Z^{\prime}}\approx\frac{\pi\alpha^{\prime 2}}{M^{2}_{\chi}}\left(1-\frac{M^{2}_{Z^{\prime}}}{M^{2}_{\chi}}\right)^{2}$ (14)
where $\alpha^{\prime}=g^{\prime 2}/4\pi$. For $g^{\prime}\sim 0.1$ and DM mass $M_{\chi}\sim 1~{}{\rm GeV}$, this leads to a cross-section which is about two orders of magnitude larger than the required ballpark to obtain correct relic density. We show the under-abundant relic density of the symmetric component in Fig. 7.
Figure 7: Under-abundant relic of symmetric component of the DM depicted by dotted red curve. The equilibrium number density is depicted by the cyan curve. Straight grey line shows the observed DM relic density.
CP-violating out of equilibrium decays $\xi_{1}^{-}\to\psi^{-}\chi,\,\xi_{1}^{+}\to(\psi^{-})^{c}\chi^{c}$ produce an asymmetry between $\chi$ and $\bar{\chi}$ as well as between $\psi$ and $\bar{\psi}$, similar to lepton asymmetry. The symmetric component of $\psi$ gets diluted by three annihilation processes: $\overline{\psi}\psi\to{\rm SM~{}SM},\,\overline{\psi}\psi\to\overline{\chi}\chi,\,\overline{\psi}\psi\to Z^{\prime}Z^{\prime}$. Note that, in presence of only the process $\overline{\psi}\psi\to{\rm SM~{}SM}$ i.e., via gauge couplings of $\psi$ to SM, $\psi$ remains under-abundant upto TeV scale, see Appendix A. If we assume that $\psi\psi Z^{\prime}$ coupling is of the same order as that of $\chi\chi Z^{\prime}$, the other two channels dominates over $\overline{\psi}\psi\to{\rm SM~{}SM}$, and the symmetric doublet component remains many orders below the correct relic ballpark for GeV scale DM. The asymmetric number density of $\psi$ gets converted to a net $\chi$ density through the decay process: $\psi\to\chi\bar{f}f$ (we assume $M_{\psi}>M_{\chi}$), induced via the soft $\mathcal{Z}_{2}$ symmetry breaking term $\mu^{2}H^{\dagger}\eta$. The symmetric part of DM $\chi$ annihilates into pairs of $Z^{\prime}$ as discussed above. Therefore, we can safely assume $\chi$ relic to be asymmetric, solely produced from CP-violating out of equilibrium decays $\xi_{1}^{-}\to\psi^{-}\chi,\,\xi_{1}^{+}\to(\psi^{-})^{c}\chi^{c}$. Feynman diagrams responsible for generation of asymmetries in dark sector are shown in Fig. 8.
Figure 8: Tree level and self energy correction diagrams producing the dark matter asymmetry.
The amount of CP-asymmetry can be quantified as,
$\epsilon_{\chi}$ $=$ $[Br(\xi_{1}^{-}\rightarrow\psi^{-}\chi)-Br(\xi_{1}^{+}\rightarrow(\psi^{-})^{c}\chi^{c})]$ (15)
$=$ $\frac{Im\left(\lambda_{1}^{*}\lambda_{2}\mathop{\sum_{k,l}}f^{*}_{1kl}f_{2kl}\right)}{8\pi^{2}(M_{2}^{2}-M_{1}^{2})}\left[\frac{M_{1}^{2}M_{2}}{\Gamma_{1}}\right]=-\epsilon_{L}.$
Figure 9: Relic density of asymmetric component of the DM obtained from the decay of $\eta$.
The abundance of asymmetric $\chi$ density can be estimated from the coupled Boltzmann equations:
$\begin{split}\frac{dY_{\xi_{1}}}{dx}&=-\frac{x}{H(M_{\xi_{1}})}s<\sigma|v|_{(\xi_{1}\xi_{1}\rightarrow All)}>\left[Y_{\xi_{1}}^{2}-{Y_{\xi_{1}}^{eq}}^{2}\right]\\ &-\frac{x}{H(M_{\xi_{1}})}\Gamma_{(\xi_{1}\rightarrow All)}\left[Y_{\xi_{1}}-Y_{\xi_{1}}^{eq}\right]\,,\\ \frac{dY_{\chi-\overline{\chi}}}{dx}&=\epsilon_{\chi}\frac{x}{H(M_{\xi_{1}})}\Gamma_{(\xi_{1}\rightarrow All)}B_{\chi}\left[Y_{\xi_{1}}-Y_{\xi_{1}}^{eq}\right].\end{split}$ (16)
Here we use $|\epsilon_{L}|=|\epsilon_{\chi}|=4.5\times 10^{-7}$. Other parameters are also kept same as that in Sec. III and Sec. IV i.e., $f=10^{-4},\,\lambda=10^{-7},\,\sigma|v|_{(\xi_{1}\xi_{1}\rightarrow All)}=10^{-25}\,{\rm GeV}^{-2}$. With this, we obtain the observed relic density for $M_{\chi}=0.70~{}{\rm GeV}$ as shown in Fig. 9. The solid red curve shows the evolution of the number density of $\xi_{1}$ particles and the solid blue line depicts the evolution of $Y^{\rm asy}_{\chi}$. Note that the observed baryon asymmetry does not get affected by the decay $\xi_{1}\to\psi\chi$, since $\psi$ and $\chi$ being vector-like fermions, do not get converted to baryon asymmetry by the sphalerons Arina:2011cu ; Arina:2012fb ; Arina:2012aj ; Narendra:2017uxl .
VI Dark Matter Self-interaction
To alleviate the small-scale anomalies of $\Lambda{\rm CDM}$, the typical DM self-scattering cross-section should be $\sigma\sim 1~{}{\rm cm}^{2}/{\rm g}\approx 2\times 10^{-24}~{}{\rm cm}^{2}/{\rm GeV}$, which is 14 orders of magnitude larger than the typical WIMP cross-section($\sigma\sim 10^{-38}\,{\rm cm}^{2}/GeV$). Besides, the self-interaction is desired to be stronger for smaller DM velocities such that it can have a large impact on small scale structures. On the other hand, it must get reduced for larger DM velocities to remain consistent with large scale $\Lambda{\rm CDM}$ predictions. SIDM with an MeV scale mediator fulfils all this requirements. $U(1)_{D}$ vector boson $Z^{\prime}$ in our model serves this purpose. The term $g^{\prime}\overline{\chi}\chi Z^{\prime}$ given in Eq. 2 enables DM self-interaction. Self-interaction can be represented by the Fynman diagram in Fig. 10.
Figure 10: Feynman diagram for DM self-interaction.
In non-relativistic limit, scattering can be described by Yukawa-type potential,
$V(r)=\pm\frac{g^{\prime 2}}{4\pi r}e^{-M_{Z^{\prime}}r}$ (17)
where, the + (-) sign denotes repulsive (attractive) potential. For vector mediators, $\chi\chi$ scattering is repulsive, which is the relevant case for this work. To capture the forward scattering divergence, we define transfer cross-section $\sigma_{T}$ as Feng:2009hw ; Tulin:2013teo ; Tulin:2017ara ,
[…]
Three distinct regimes can be identified for the scattering.
  • Born regime $\left(g^{\prime 2}M_{\chi}/(4\pi M_{Z^{\prime}})\ll 1,\frac{M_{\chi}v}{M_{Z^{\prime}}}\geq 1\right)$
  • Classical regime $\left(g^{\prime 2}M_{\chi}/(4\pi M_{Z^{\prime}})\geq 1,\frac{M_{\chi}v}{M_{Z^{\prime}}}\geq 1\right)$
  • Resonance Regime $\left(g^{\prime 2}M_{\chi}/(4\pi M_{Z^{\prime}})\geq 1,\frac{M_{\chi}v}{M_{Z^{\prime}}}\leq 1\right)$
Cross-sections for each regime are listed in Appendix B. Perturbative calculations hold good only in the Born regime. In the resonant regime, for attractive potential we see quantum mechanical resonances for smaller velocities due to (quasi)-bound state formation Dutta:2022knf . Since in the present case, we have asymmetric DM with an repulsive potential, we do not see any resonance spikes. In Fig. 11, we show the self-interaction allowed parameter space for $g^{\prime}=0.1$ in $M_{\chi}$ $M_{Z^{\prime}}$ plane. We constrain $\sigma/M_{\chi}$ as:
  • $0.1-10~{}{\rm cm}^{2}/{\rm g}$ for dwarfs ($v\sim 10~{}\rm km/s$)
  • $0.1-1~{}{\rm cm}^{2}/{\rm g}$ for galaxies ($v\sim 100~{}\rm km/s$)
  • $0.1-1~{}{\rm cm}^{2}/{\rm g}$ for clusters ($v\sim 1000~{}\rm km/s$)
These regions are depicted by cyan, green and magenta colored regions in Fig. 11 as indicated in figure inset.
Figure 11: Parameter space for self-interaction in $M_{\chi}$ $M_{Z^{\prime}}$ plane with $g^{\prime}=0.1$.
Figure 12: Self-interaction cross-section per unit mass of DM in terms of average collision velocity.
The top corner corresponds to the Classical regime while the bottom corner corresponds to the Born regime, with resonant regime sandwiched between the two regimes.
We also plot self-scattering cross-section per unit DM mass versus average collision velocity in Fig. 12 for different DM masses. Keeping $M_{Z^{\prime}}=0.008~{}{\rm GeV}$ and $g^{\prime}=0.1$, we plot the cross-section per unit DM mass for $M_{\chi}=0.7~{}{\rm GeV},1~{}{\rm GeV},5~{}{\rm GeV}$. It fits data from dwarfs (red), low surface brightness (LSB) galaxies (blue), and clusters (green) Kaplinghat:2015aga ; Kamada:2020buc , for the values of $M_{\chi}$, $M_{Z^{\prime}}$ as indicated in the figure. It is clear that the model can appreciably explain velocity-dependent DM self-interaction.
VII DM Direct Detection
DM $\chi$ undergoes elastic scattering off detector nucleons. This is facilitated by $Z-Z^{\prime}$ mixing as shown in Fig. 13.
Figure 13: Elastic scattering of DM $\chi$ off a nucleon mediated by $Z-Z^{\prime}$ mixing.
Let us consider spin independent direct search, the cross-section for which is given by,
$\sigma^{\rm SI}_{\chi N}=\frac{g^{2}g^{\prime 2}\epsilon^{2}}{\pi}\frac{\mu^{2}_{\chi N}}{M^{4}_{Z^{\prime}}}\frac{\big{(}Zf_{p}+(A-Z)f_{n}\big{)}^{2}}{A^{2}}.$ (19)
where, $\mu_{\chi N}=\frac{M_{\chi}m_{N}}{(M_{\chi}+m_{N})}$ is the reduced mass of the DM-nucleus ($N$) system with proton number $Z$ and mass number $A$. $\epsilon$ is the $Z-Z^{\prime}$ kinetic mixing, $f_{p}$ ($f_{n}$) are interaction strengths for proton (neutron) respectively.
We can constrain model parameters with data from direct search experiments like CRESST-III Abdelhameed:2019hmk and XENON1T Aprile:2018dbl . In Fig. 14, the most stringent constraints from CRESST-III (dotted contours in black) and XENON1T (dotted contours in pink) experiments are shown assuming $g^{\prime}=0.1$. For below (above) $10{\rm GeV}$ DM, we use constraints from CRESST-III (XENON1T) experiment. The region towards left of each contour is excluded for the mixing parameter mentioned alongside the curves.
Figure 14: Constraints from DM direct detection in the plane of DM mass $(M_{\chi})$ versus mediator mass $(M_{Z^{\prime}})$ with $g^{\prime}=0.1$.
VIII Conclusions
Neutrino oscillation experiments convincingly establishes non-zero neutrino mass. However, no consensus has been established regarding Dirac or Majorana nature for neutrinos, as experiments looking for neutrinoless double beta decay does not provide any conclusive evidence. In this work, we assume neutrinos to be Dirac particles. Then we explain baryon and dark matter asymmetry in a common framework.
We extend SM with a $SU(2)_{L}$ doublet $\psi$ and a singlet $\chi$. The singlet fermion $\chi$ being odd under the discrete $\mathcal{Z}_{2}$ symmetry behave as a candidate of DM. The same $\mathcal{Z}_{2}$ symmetry disallowed neutrino Dirac mass. $\mathcal{Z}_{2}$ symmetry is softly broken without destabilizing DM $\chi$, ensured by $M_{\chi}<M_{\psi}$. This also leads to a Dirac mass of neutrinos. A heavy scalar doublets ($\eta$), having non-trivial $\mathcal{Z}_{2}$ charge, undergoes out-of-equilibrium decay. Decay channels $\eta\to\nu_{R}\ell$ and $\eta\to\chi\psi$ generates lepton and dark matter asymmetries. As $B-L$ becomes an exact symmetry, the CP-violating decay of $\eta$ to $\nu_{R}\ell$ produces equal and opposite $B-L$ asymmetries in the left and right-handed sectors. The right-handed sector is weakly coupled to SM sector, required by neutrino Dirac mass. $B-L$ asymmetry in the left-handed sector converts to a net $B$-asymmetry via the $B+L$ violating sphaleron transitions. The asymmetries in the right-handed sector remains intact. The asymmetries neutralizes much after the electroweak phase transition when the sphaleron transitions gets suppressed. Meanwhile the symmetric DM component annihilates away to pairs of light mediators. The light mediators facilitates self-interaction among the DM particles, a essential feature for alleviating the small-scale $\Lambda{\rm CDM}$ anomalies. The kinetic mixing between dark and visible sectors facilitates direct detection of DM. We have found the relevant parameter space for SIDM confronted with direct search constraints from experiments like CRESST-III and XENON-1T. Since kinetic mixing suppresses all annihilations to SM final states, all gamma, cosmic, and neutrino fluxes are well below the near future reach of indirect probes Arina:2018zcq ; https://doi.org/10.1002/andp.19163561802 ; Cassel:2009wt ; Iengo:2009ni ; Slatyer:2009vg . The indirect signals of SIDM remains much below the current constraints by the experiments like Fermi-LAT Fermi-LAT:2013sme ; Fermi-LAT:2015att , MAGIC MAGIC:2016xys , HESS HESS:2018cbt , AMS-02 PhysRevLett.110.141102 , CMB by Planck Aghanim:2018eyx , and $\gamma$-rays by INTEGRAL Knodlseder:2007kh . Although $\chi$ is invisible at the collider, the doublet component $\psi^{\pm}$ can be pair produced in the collider  Bhattacharya:2015qpa ; Bhattacharya:2017sml , and their decay may lead to interesting signatures.
Appendix A Relic Abundance of Doublet DM having only gauge interactions with the SM
If the DM is a vector-like fermion doublet $\psi$ of the form $\psi^{T}=(\psi^{0},\psi^{-})$ (with hypercharge $Y=-1$, where we use $Q=T_{3}+Y/2$), odd under an imposed $\mathcal{Z}_{2}$ symmetry, its neutral component becomes stable DM. The Lagrangian for the doublet DM is given by,
$\mathcal{L}=\overline{\psi}\left(i\gamma^{\mu}D_{\mu}-M_{\psi}\right)\psi$ (20)
where, the covariant derivative involving the $SU(2)_{L}$ gauge boson triplet $W_{\mu}^{a}$  $(a=1,2,3)$ and $U(1)_{Y}$ gauge boson $B_{\mu}$ given by:
$D_{\mu}=\partial_{\mu}-i\frac{g}{2}\tau_{a}.W_{\mu}^{a}-ig_{Y}\frac{Y}{2}B_{\mu}$ (21)
where $\tau_{a}$ are Pauli spin matrices (generators of $SU(2)$), $g$ and $g_{Y}$ denote $SU(2)$ and $U(1)$ coupling strengths respectively. Expanding the covariant derivative of the Lagrangian given in Eq.20, we get the interaction term of $\psi^{0}$ and $\psi^{\pm}$ with the SM gauge bosons as follows:
$\mathcal{L}_{int}$ $=\overline{\psi}i\gamma^{\mu}(-i\frac{g}{2}\tau.W_{\mu}-ig_{Y}\frac{Y}{2}B_{\mu})\psi$ (22)
$=\Big{(}\frac{e}{2\sin\theta_{W}\cos\theta_{W}}\Big{)}\overline{\psi^{0}}\gamma^{\mu}Z_{\mu}\psi^{0}$
$+\frac{e}{\sqrt{2}\sin\theta_{W}}(\overline{\psi^{0}}\gamma^{\mu}W^{+}_{\mu}\psi^{-}+\psi^{+}\gamma^{\mu}W^{-}_{\mu}\psi^{0})$
$-e\psi^{+}\gamma^{\mu}A_{\mu}\psi^{-}$
$-\Big{(}\frac{e\cos 2\theta_{W}}{2\sin\theta_{W}\cos\theta_{W}}\Big{)}~{}\psi^{+}\gamma^{\mu}Z_{\mu}\psi^{-}.$
where $g=\frac{e}{\sin\theta_{W}}$ and $g_{Y}=\frac{e}{\cos\theta_{W}}$, $e$ and $\theta_{W}$ are the elementary charge and Weinberg angle respectively. All interactions being fixed by the SM gauge charge, the model does not have any free parameter except the mass $M_{\psi}$. So both relic density and direct search cross-sections become a function of DM mass alone. The model has been implemented in LanHEP SEMENOV1998124 and MicrOmegas Belanger:2013oya . The relic density obtained from the model as a function of DM mass is shown in Fig. 15. We observe that correct relic density is obtained only for DM mass around 1 TeV.
Figure 15: Relic density of doublet DM with only gauge interaction with SM.
Appendix B DM Self-interaction Cross-sections at Low Energy
In the Born Limit ($\alpha^{\prime}M_{\chi}/M_{Z^{\prime}}<<1$), for both attractive as well as repulsive potentials, the transfer cross-section is
$\sigma^{\rm Born}_{T}=\frac{8\pi\alpha^{\prime 2}}{M^{2}_{\chi}v^{4}}\Bigg{(}\ln(1+M^{2}_{\chi}v^{2}/M^{2}_{Z^{\prime}})-\frac{M^{2}_{\chi}v^{2}}{M^{2}_{Z^{\prime}}+M^{2}_{\chi}v^{2}}\Bigg{)}.$ (23)
Outside the Born regime ($\alpha^{\prime}M_{\chi}/M_{Z^{\prime}}\geq 1$), we have two distinct regions viz the classical region and the resonance region. In the classical limit ($M_{\chi}v/M_{Z^{\prime}}\geq 1$), the solutions for an attractive potential is given by Tulin:2013teo ; Tulin:2012wi ; Khrapak:2003kjw
$\sigma^{\rm classical}_{T}(\rm Attractive)=\left\{\begin{array}[]{l}\frac{4\pi}{M^{2}_{Z^{\prime}}}\beta^{2}\ln(1+\beta^{-1});~{}~{}\beta\leqslant 10^{-1}\\ \frac{8\pi}{M^{2}_{Z^{\prime}}}\beta^{2}/(1+1.5\beta^{1.65});10^{-1}\leq\beta\leqslant 10^{3}\\ \frac{\pi}{M^{2}_{Z^{\prime}}}(\ln\beta+1-\frac{1}{2}ln^{-1}\beta);~{}~{}\beta\geq 10^{3}\\ \end{array}\right.$ (24)
In the classical limit ($M_{\chi}v/M_{Z^{\prime}}\geq 1$), the solutions for repulsive potential is given byTulin:2013teo ; Tulin:2012wi ; Khrapak:2003kjw
$\sigma^{\rm classical}_{T}(\rm Repulsive)=\left\{\begin{array}[]{l}\frac{2\pi}{M^{2}_{Z^{\prime}}}\beta^{2}\ln(1+\beta^{-2});~{}~{}\beta\leqslant 1\\ \frac{\pi}{M^{2}_{Z^{\prime}}}(\ln 2\beta^{2}-\ln~{}\ln 2\beta)^{2};~{}~{}\beta\geq 1\\ \end{array}\right.$ (25)
where $\beta=2\alpha^{\prime}M_{Z^{\prime}}/(M_{\chi}v^{2})$.
In the resonant regime ($g^{\prime 2}M_{\chi}/(4\pi M_{Z^{\prime}})\geq 1,M_{\chi}v/M_{Z^{\prime}}\leq 1$), the quantum mechanical resonances and anti-resonance in $\sigma_{T}$ appear due to (quasi-)bound states formation in the attractive potential. In the resonant regime, non-perturbative results obtained by approximating the Yukawa potential to be a Hulthen potential $\Big{(}V(r)=\pm\frac{g^{\prime 2}}{4\pi}\frac{\delta e^{-\delta r}}{1-e^{-\delta r}}\Big{)}$ is given by Tulin:2013teo :
$\sigma^{\rm Hulthen}_{T}=\frac{16\pi\sin^{2}\delta_{0}}{M^{2}_{\chi}v^{2}}$ (26)
where $l=0$ phase shift is given in terms of the $\Gamma$ functions by
$\delta_{0}={\rm arg}\Bigg{(}\frac{i\Gamma\Big{(}\frac{iM_{\chi}v}{kM_{Z^{\prime}}}\Big{)}}{\Gamma(\lambda_{+})\Gamma(\lambda_{-})}\Bigg{)}$ (27)
where,
$\lambda_{\pm}=\left\{\begin{array}[]{l}1+\frac{iM_{\chi}v}{2kM_{Z^{\prime}}}\pm\sqrt{\frac{\alpha_{D}M_{\chi}}{kM_{Z^{\prime}}}-\frac{M^{2}_{\chi}v^{2}}{4k^{2}M^{2}_{Z^{\prime}}}}~{}~{}~{}~{}{\rm Attractive}\\ 1+\frac{iM_{\chi}v}{2kM_{Z^{\prime}}}\pm i\sqrt{\frac{\alpha_{D}M_{\chi}}{kM_{Z^{\prime}}}+\frac{M^{2}_{\chi}v^{2}}{4k^{2}M^{2}_{Z^{\prime}}}}~{}~{}~{}~{}{\rm Repulsive}\\ \end{array}\right.$ (28)
with $k\approx 1.6$ being a dimensionless number.